OpenStudy (anonymous):

A right cone has a volume of 471 and a height of 18. What is the radius?

OpenStudy (goldphenoix):

\[V=Pir^2h/3\] \[471=(3.14*r^2*18)/3\] \[471=56.52r^2/3\] \[471*3=56.52r^2\] \[1413=56.52r^2\] \[1413/56.52=r^2\] \[25=r^2\] R=5 Sorry. This isn't my best work. Really tired. I'm going to sleep. I hope this help!

OpenStudy (anonymous):

It helps a lot, Thank you.